# RoDUS: Robust Decomposition of Static and Dynamic Elements in Urban Scenes

### v1

@@ Abstract
The task of separating dynamic objects from static environments using NeRFs has been widely studied in recent years. However, capturing large-scale scenes still poses a challenge due to their complex geometric structures and unconstrained dynamics. Without the help of 3D motion cues, previous methods often require simplified setups with slow camera motion and only a few/single dynamic actors, leading to suboptimal solutions in most urban setups. To overcome such limitations, we present RoDUS, a pipeline for decomposing static and dynamic elements in urban scenes, with thoughtfully separated NeRF models for moving and non-moving components. Our approach utilizes a robust kernel-based initialization coupled with 4D semantic information to selectively guide the learning process. This strategy enables accurate capturing of the dynamics in the scene, resulting in reduced artifacts caused by NeRF on background reconstruction, all by using self-supervision. Notably, experimental evaluations on KITTI-360 and Pandaset datasets demonstrate the effectiveness of our method in decomposing challenging urban scenes into precise static and dynamic components.
@@ 1 Introduction
There have been numerous works that achieved success in various 3D representations including both implicit [23] and explicit [36, 37] scene modeling. Therefore, reconstructing and decomposing large-scale dynamic outdoor scenes is becoming a critical task for 3D scene understanding, especially for domains such as closed-loop simulation for autonomous driving [48, 54, 40]. Accurate segmentation and removal of dynamic objects are crucial as they cause floating artifacts during novel view synthesis (NVS) of background scenes due to view inconsistency. This inevitably affects other downstream applications. For instance, several works [20, 7, 12] require masking out eventual dynamic objects as they estimate camera poses. Indeed, in the case of highly dynamic scenes, without feeding the correct masks, they tend to make wrong refinements of the poses. On the other hand, dynamic objects often contain valuable information in the scene and play a main role in tasks such as surveillance and video compression [22].
@@
However, advancements are lagging as understanding time-dependent scenes is still an ill-posed problem. This is because it demands considerable efforts to detect and treat moving components separately, while still ensuring spatio-temporal consistency compared to static scenes [30, 44]. The challenges further arise in driving settings, featuring not only large unbounded scenes, with lots of unconstrained dynamics and various illuminations but also large spatio-temporal gaps captured from the ego-vehicle. This illustrates that a majority of the background regions are captured only for a short duration, due to high occlusion caused by other dynamic objects while the ego-vehicle is also in continuous motion. Unlike previous works [28, 29, 18, 47], where scenes typically involve a single object with slow ego-camera motion, the scenarios encountered in driving sequences are less controllable.
@@
Until now, the research community has proposed various solutions to the challenge of decoupling dynamic objects from static environments. Many of those develop NeRF-based multi-branch architectures in which only the dynamic branch (i.e. time-dependent branch) is employed for rendering dynamic objects while the background is learned using a separate static branch. Subsequently, these methods need to detect moving objects offline, so that the static model could later avoid sampling points at regions that belong to dynamic objects [27, 15, 48, 50, 53, 60, 51]. Nevertheless, they end up relying on expensive annotations for segmenting and tracking every object on the scene to enable correct model guidance. Such limitation restricts their adaptability to other datasets. Others [21, 4] learn to classify the inconsistency through a rejection factor while jointly optimizing both branches. However, NeRF’s view-dependent capacity makes it difficult to distinguish non-Lambertian but static from true dynamic objects, requiring careful and expensive parameter tuning to reach optimal separation.
@@
In this paper, by building upon self-supervised 4D scene representation using NeRF, RoDUS (Robust Decomposition of Static and Dynamic Elements in Urban Scenes) employs a two-pathway architecture with additional sky and shadow modeling that is suitable for urban scenes. While the architecture is mainly used for representing dynamic scenes as a whole, we argue that there appears the possibility of both branches learning incorrect information (i.e. local minima), as long as the overall reconstruction satisfies the ground truth image. This deviates from the initial design intent, where the static branch is meant to learn only the static background, and vice versa. Toward this goal, we inspire from [1] by proposing a robust kernel-based strategy, thus reducing L2 sensitivity. Additionally, we introduce a view-independent semantic field to prevent the static model from expressing unwanted regions, while still being able to refine high-frequency detail. This semantic awareness is further used to guide the information flow, thus balancing the contribution of each branch. To summarize, the main contributions of our work are:
@@
  • We conduct comprehensive comparisons on several challenging driving sequences, demonstrating our RoDUS’s performance on static-dynamic components extraction, and decoupling compared to other SOTA methods.
@@ 2 Related Work
Recently, many works have achieved promising results for high-fidelity view synthesis [2], increasing the training and inference speeds [25], and combining with semantic [9, 15, 58, 13], foundation model-guided [14, 41] reasoning to suit urban driving scenarios [33, 10]. As most of the above methods focus on static scenes, some other works have targeted non-static, transient elements from video sequences [28, 29, 18, 17, 30, 44].
NeRFs for outliers rejection. NeRF-W [21] is the pioneering work that decomposes the scene into static and transient components to reconstruct the landmark from unconstrained phototourism datasets. They propose per-frame latent embeddings and a transient head to model different lighting conditions and transient effects, thus these phenomena can be removed from the rendering function at inference. Alternatively, RobustNeRF [34] successfully rejects distractors from the scenes using a trimmed kernel to modify the photometric loss into an Iteratively Least Squares problem, they also propose patch filtering to maintain sharp reconstruction quality. Ha-NeRF [4], instead, parameterizes the kernel as learnable and uses the 2D visibility map per image as a weight function. This occlusion-aware module is proven to be more accurate at separating transient subjects.
Motion segmentation. In addition to advancements in 3D vision, there has been extensive research focusing on self-supervised motion segmentation at 2D image level. The majority of these methods rely on 2D trackers [7], optical flow patterns [3, 16, 55] and/or image warping [35] to distinguish between ego motions and object motions and segment them. These approaches have clear limitations as they focus solely on image-level processing and lack scene-level understanding. Therefore, they struggle to handle complicated static backgrounds with large camera motion [47]. Furthermore, those who follow the pipeline of first-process-then-NeRF create not only label ambiguity that affects learning but also add extra costs for separate perception processing tasks, considering that motion segmentation or flow classification themselves are already challenging.
@@
SUDS [43], and EmerNeRF [52] are the most similar to our approach since they utilize 2D off-the-shelf models to enable object- and scene-level semantic understanding while still being able to separate dynamic objects. Our approach, however, differs from existing works: we pay more attention to the representation of each branch individually, a factor often neglected by other studies but stands advantageous in handling such challenging data. This gives us better performance in lots of downstream tasks such as background reconstruction, object removal, and motion segmentation.
@@ 3 Method
This section details RoDUS’s methodology of disentangling and composing representation in addition to its computational pipeline which is illustrated in Fig. 2. We describe in Sec. 3.1 the architecture for our scene representation. In Sec. 3.2, we introduce the semantic-radiance field used in RoDUS, how it can be rendered using differentiable volume rendering and contributes to the separation. Finally, Sec. 3.3 formulates the losses to train RoDUS as well as our proposed training strategy, which is demonstrated in Sec. 3.4.
Goals. Our main goal is to learn a scene representation of a dynamic environment (e.g. driving scene) with the ability to semantically disentangle between static components (e.g. road, building, vegetation, parked vehicles) and dynamic objects in a self-supervised manner. RoDUS takes a set of $N$ input RGB images of size $W\times H$, their associated poses, and capturing timestamps $\{I_{i}^{H\times W\times 3},p_{i},t_{i}\}_{i=1}^{N}$. Additionally, our model relies on semantic $\{S_{i}^{H\times W\times L}\}_{i=1}^{N}$, where $L$ is the number of classes, and geometric cues obtained from pre-trained 2D segmentation model and LiDAR depth measurements, respectively. At inference time, we wish to achieve rendered images for static $\{\hat{I}_{i}^{S},\hat{S}_{i}^{S}\}_{i=1}^{N}$ and dynamic $\{\hat{I}_{i}^{D},\hat{S}_{i}^{D}\}_{i=1}^{N}$ components of the scene for each timestamp independently, which can also be combined for a complete scene. As mentioned above, we study the quality of individual components rather than conforming solely to a satisfying composed result.
@@ 3.1 Scene Representation
Preliminaries. RoDUS is based on NeRF [23] which is designed to capture both geometry and view-dependent appearance of a scene from a set of RGB images. It encodes the scene within a Multi-Layer Perceptron (MLP) and outputs color $\mathbf{c}(\mathbf{x},\mathbf{d})\in\mathbb{R}^{3}$ and density $\sigma(\mathbf{x})\in\mathbb{R}$ at any query position $\mathbf{x}\in\mathbb{R}^{3}$ along with its viewing direction $\mathbf{d}\in\mathbb{R}^{3}$. During training, using intrinsic parameters and poses, every pixel of the image undergoes a ray marching procedure $\mathbf{r}$ to sample inputs for the MLP. The model parameters are optimized to predict the colors rendered along the ray using the mean squared error: $\mathcal{L}=\|\hat{\mathbf{C}}_{i}(\mathbf{r})-\mathbf{C}_{i}(\mathbf{r})\|_{2}^{2}$.
Architecture. Inspired by [21, 18, 47], we factorize the scene into two elements, each learned through a separate pathway: one for moving components and the other for non-moving background. Each branch is modeled independently using its own multi-level hash grid representation and MLPs. This approach enables the concurrent learning of both static and dynamic aspects of the scene.
The static branch whose mission is to learn all non-moving parts in the scene, uses a 3D hash grid $\mathcal{H}^{S}$, analogous to the one used in [25]. To model lighting variations, we also condition on a per-frame latent embedding $l_{a}$ [21]:
@@
As the dynamic branch allows its outputs to depend on time, we employ a 4D hash grid $\mathcal{H}^{D}$ with one additional normalized timestamp $t$ as input. The dynamic field additionally incorporates a shadow head [47] designed to learn a shadow ratio $\rho\in[0,1]$. This scalar is later used to scale down the static radiance values $\mathbf{c}^{S}$ in Eq. 3:
@@
Figure 2: RoDUS’s architecture. Our model comprises two separate branches that take as input sampled positions $\mathbf{x}$, their viewing direction $\mathbf{d}$, and their timestamp $t$ for the dynamic branch, and generate outputs for every query coordinate. Each branch is represented by a separate hash grid and the MLP-based neural function, which predicts colors, densities, and semantics. The rendered static image is used to calculate IRLS map $\omega(\epsilon)$ during the robust initialization step (Sec. 3.3), while dynamic semantic outputs are passed through a “foreground-only mask” to prevent over-explaining background regions (Sec. 3.2).
Rendering. We calculate the color $\hat{C}$ for each camera ray $\mathbf{r}$ by integrating all the sampled points along the ray:
@@
where $T_{i}=\exp\left(-\sum_{j=1}^{i-1}(\sigma_{j}^{S}+\sigma_{j}^{D})\delta_{j}\right)$ is the accumulated transmittance and $\alpha_{i}^{\xi}=1-\exp(-\sigma_{i}^{\xi}\delta_{i})$ with $\xi\in\{S,D\}$ are alpha values. Given the outputs of each branch, we can render static and dynamic maps separately.
@@ 3.2 Enabling Semantic Awareness
NeRF provides low-level implicit representations of geometry and radiance but lacks a higher-level understanding of the scene. It is shown in [34] that relying only on view-dependent radiance and sensitive mean square error makes the model prone to outliers and over-fit observations. On the other hand, semantic labels can be conceptualized as a view-invariant function [59], mapping world coordinates $\mathbf{x}$ to a distribution over semantic labels $\mathbf{s}(\mathbf{x})\in\mathbb{R}^{L}$. As a result, we choose to integrate a semantic head to predict the semantic class for each point, similar to the color head. Subsequently, the same rendering formula is applied for semantics to achieve 2D semantic segmentation maps:
@@
Figure 3: (a) We enforce a “foreground-only mask” to the dynamic semantic head. In the forward pass (black arrows), the mask prevents the dynamic branch from outputting pixels that do not belong to foreground classes. Similarly, in the backward pass (orange arrows), it restricts the dynamic field from learning background pixels, which may primarily come from noisy annotations. (b) Predictions generated by 2D segmentation model are noisy and inconsistent between views. (c) Since the dynamic field includes a temporal dimension, it ends up learning these conflicting labels, satisfying the overall loss. (d) Therefore, our proposed mask is used to tackle the problem.
@@
We apply softmax normalization at each query point individually before ray integration, as introduced in [39]. This is done due to the unbounded nature of logits, ensuring that the semantic head can not generate excessively high scores in regions with low density to overcome high-density areas. Remarkably, this not only improves the stability of rendering semantics but also prevents half-precision floating point (float16) overflowing when working with frameworks such as tiny-cuda-nn [24].
(a) Ground-truth

### v2

@@ Abstract
The task of separating dynamic objects from static environments using NeRFs has been widely studied in recent years. However, capturing large-scale scenes still poses a challenge due to their complex geometric structures and unconstrained dynamics. Without the help of 3D motion cues, previous methods often require simplified setups with slow camera motion and only a few/single dynamic actors, leading to suboptimal solutions in most urban setups. To overcome such limitations, we present RoDUS, a pipeline for decomposing static and dynamic elements in urban scenes, with thoughtfully separated NeRF models for moving and non-moving components. Our approach utilizes a robust kernel-based initialization coupled with 4D semantic information to selectively guide the learning process. This strategy enables accurate capturing of the dynamics in the scene, resulting in reduced floating artifacts in the reconstructed background, all by using self-supervision. Notably, experimental evaluations on KITTI-360 and Pandaset datasets demonstrate the effectiveness of our method in decomposing challenging urban scenes into precise static and dynamic components.
@@ 1 Introduction
There have been numerous works that achieved success in various 3D representations including both implicit [23] and explicit [36, 37] scene modeling. Therefore, reconstructing and decomposing large-scale dynamic outdoor scenes is becoming a critical task for 3D scene understanding, especially for domains such as closed-loop simulation for autonomous driving [48, 54, 40]. Accurate segmentation and removal of dynamic objects are crucial as they are the main reason for artifacts during novel view synthesis (NVS) of background regions due to view inconsistency. This inevitably affects other downstream applications. For instance, several works [20, 7, 12] require masking out eventual dynamic objects as they estimate camera poses. Indeed, in highly dynamic scenes, without feeding the correct masks, they tend to make wrong refinements of the poses. On the other hand, dynamic objects often contain valuable information in the scene and play a main role in tasks such as surveillance and video compression [22].
@@
However, advancements are lagging as understanding time-dependent scenes is an ill-posed problem. This is because it demands considerable efforts to detect and treat moving components separately while ensuring spatiotemporal consistency compared to static scenes [30, 44]. The challenges further arise in driving settings, featuring not only large unbounded scenes, with lots of unconstrained dynamics and various illuminations but also large temporal gaps captured from the ego-vehicle. This demonstrates that a majority of the background regions are captured only for a short duration, due to high occlusions caused by other dynamic objects while the ego-vehicle is also in continuous motion. Unlike previous works [28, 29, 18, 47], where scenes typically involve a single object with slow ego-motion, the scenarios encountered in driving sequences are less controllable.
@@
Until now, the research community has proposed various solutions to the challenge of decoupling dynamic objects from static environments. Many develop NeRF-based multi-branch architectures in which only the dynamic (i.e. time-dependent) model is employed for rendering dynamic objects while the background is learned using a separate static model. Subsequently, these methods need to detect moving objects offline, so that the static model could later avoid sampling points at regions that belong to dynamic objects [27, 15, 48, 50, 53, 60, 51]. Nevertheless, they end up relying on expensive annotations for segmenting and tracking every object in the scene to enable correct guidance. Such limitation restricts their adaptability to many datasets. Others [21, 4] learn to classify the inconsistency through a rejection factor while jointly optimizing both branches. However, NeRF’s view-dependent capacity makes it difficult to distinguish non-Lambertian but static from true dynamic objects, requiring careful and expensive parameter tuning to reach optimal separation.
@@
In this paper, by building upon self-supervised 4D scene representation using NeRF, RoDUS (Robust Decomposition of Static and Dynamic Elements in Urban Scenes) employs a two-pathway architecture with additional sky, road, and shadow modeling that is suitable for urban scenes. While the architecture has been mainly used for representing dynamic scenes as a whole, we argue that it appears a possibility of both branches learning incorrect information (i.e. local minima), as long as the overall reconstruction satisfies the ground truth image. This deviates from the initial design intent, where the static model is meant to learn only the static background, and vice versa. Toward this goal, we draw inspiration from [1] by proposing a robust kernel-based strategy, thus reducing L2 sensitivity on the static model. Additionally, we introduce a view-independent semantic field to prevent the dynamic model from expressing unwanted regions, while still being able to refine high-frequency detail. This semantic awareness is further used to guide the information flow, thus balancing the contribution of each branch. To summarize, the main contributions of our work are:
@@
  • We conduct comprehensive comparisons on several challenging driving sequences, indicating our RoDUS’s performance on static-dynamic components extraction, and decoupling compared to other SoTA methods.
@@ 2 Related Work
Recently, many works have achieved promising results in high-fidelity view synthesis [2] and combining with semantic [9, 15, 58, 13], foundation model-guided [14, 41] reasoning to better accommodate urban driving scenarios [33, 10]. As most of the above methods focus on static scenes, some other works have targeted non-static, transient elements from video sequences [28, 29, 18, 17, 30, 44].
NeRFs for outliers rejection. NeRF-W [21] is the pioneering work that decomposes the scene into static and transient components to reconstruct landmarks from unconstrained phototourism datasets. They propose per-frame latent embeddings and a transient head to model different lighting conditions and transient effects, thus these phenomena can be removed from the rendering function at inference. Alternatively, RobustNeRF [34] successfully rejects distractors from the scenes using a trimmed kernel to modify the photometric loss into an Iteratively Least Squares problem, they also include patch filtering to maintain sharp reconstruction quality. Ha-NeRF [4], instead, parameterizes the kernel as learnable and uses 2D visibility map per image as a weight function. This occlusion-aware module is proven to be more accurate at separating transients.
Motion segmentation. In addition to advancements in 3D vision, there has been extensive research focusing on self-supervised motion segmentation at 2D image level. The majority of these methods rely on 2D trackers [7], optical flow patterns [3, 16, 55] and/or image warping [35] to distinguish between ego motions and object motions and segment them. These approaches have clear limitations as they focus solely on image-level processing and lack scene-level understanding. Therefore, they struggle to handle complicated scenes with large camera motion [47]. Furthermore, those who follow the pipeline of first-process-then-NeRF create not only label ambiguity that affects learning but also add extra costs for separate perception processing tasks, considering that motion segmentation or flow classification themselves are already challenging.
@@
SUDS [43] and EmerNeRF [52] are the most similar to our approach since they utilize 2D off-the-shelf models to enable object-/scene-level semantic understanding while still being able to separate dynamic objects. However, our approach differs from existing works in that we place greater emphasis on the representation of each branch individually, a factor often overlooked by other studies. This focus proves advantageous in understanding complex scenes. Specifically, our dynamic model is designed to learn only the true dynamic instances rather than capturing variations in illumination or pixel changes caused by ego-motion.
@@ 3 Method
This section details RoDUS’s methodology of disentangling and composing representation with its computational pipeline illustrated in Fig. 2. Sec. 3.1 describes the architecture for our scene representation. Sec. 3.2 introduces the semantic-radiance field used in RoDUS, how it can be rendered using differentiable volume rendering and contributes to the separation. Finally, we formulate the losses to train RoDUS in Sec. 3.3 as well as our proposed training strategy in Sec. 3.4.
Goals. Our main goal is to learn a scene representation of a dynamic environment (e.g. a driving scene) with the ability to semantically disentangle between static background (e.g. road, building, vegetation, parked vehicles) and dynamic objects in a self-supervised manner. RoDUS takes a set of $N$ input RGB images, their associated poses, and capturing timestamps $\{I_{i},p_{i},t_{i}\}_{i=1}^{N}$. Additionally, our model relies on semantic $\{S_{i}\}_{i=1}^{N}$, and geometric cues obtained from pre-trained 2D segmentation model and LiDAR depth measurements. At inference, we wish to achieve rendered images for static $\{\hat{I}_{i}^{S},\hat{S}_{i}^{S}\}_{i=1}^{N}$ and dynamic $\{\hat{I}_{i}^{D},\hat{S}_{i}^{D}\}_{i=1}^{N}$ components of the scene for each timestamp independently, which can also be combined for a complete view. As mentioned above, we study the quality of individual components rather than conforming solely to a satisfying composed result.
@@ 3.1 Scene Representation
Preliminaries. RoDUS is based on NeRF [23], which is designed to capture both geometry and view-dependent appearance of a scene from a set of RGB images. It encodes the scene within a Multi-Layer Perceptron (MLP) and outputs color $\mathbf{c}(\mathbf{x},\mathbf{d})\in\mathbb{R}^{3}$ and density $\sigma(\mathbf{x})\in\mathbb{R}$ at any query position $\mathbf{x}\in\mathbb{R}^{3}$ along with its viewing direction $\mathbf{d}\in\mathbb{R}^{3}$. During training, using intrinsic parameters and poses, every pixel of the image undergoes a ray marching procedure $\mathbf{r}$ to sample inputs for the MLP. The model parameters are optimized to predict the colors rendered along the ray using the mean squared error: $\mathcal{L}=\|\hat{\mathbf{C}}_{i}(\mathbf{r})-\mathbf{C}_{i}(\mathbf{r})\|_{2}^{2}$.
Architecture. Inspired by [21, 18, 47], we factorize the scene into two elements, each learned through a separate pathway. Each branch is modeled independently using its own multi-level hash grid representation and MLPs. This approach enables the concurrent learning of both aspects of the scene.
The static branch uses a 3D hash grid $\mathcal{H}^{S}$ for positions and positional embedding $\gamma(\cdot)$ for viewing directions, analogous to the one used in [25]. To model lighting variations, we also condition on a per-frame latent embedding $l_{a}$ [21]:
@@
As the dynamic branch allows its outputs to depend on time, we extend to 4D hash grid $\mathcal{H}^{D}$ with one additional normalized timestamp $t\in[0,1]$ as input. We additionally incorporates a shadow head [47], designed to learn a shadow ratio $\rho\in[0,1]$. This scalar is later used to scale down the static radiance values $\mathbf{c}^{S}$ in Eq. 3:
@@
Figure 2: RoDUS’s architecture. Our model comprises two separate branches that take as input sampled positions $\mathbf{x}$, viewing direction $\mathbf{d}$, and their timestamp $t$, and generate outputs for every query coordinate. Each branch is represented by a separate hash grid and the MLP-based neural function, which predicts colors, densities, and semantics. The rendered static RGB is used to calculate IRLS map $\omega(\epsilon)$ during the robust initialization step (Sec. 3.3), while dynamic semantic outputs are passed through a “foreground-only mask” to prevent over-explaining background regions (Sec. 3.2).
Rendering. We calculate the color $\hat{C}(\mathbf{r})$ of each ray $\mathbf{r}$ by integrating all the sampled points along the ray:
@@
where $T_{i}=\exp\left(-\sum_{j=1}^{i-1}(\sigma_{j}^{S}+\sigma_{j}^{D})\delta_{j}\right)$ is the accumulated transmittance and $\alpha_{i}^{\xi}=1-\exp(-\sigma_{i}^{\xi}\delta_{i})$ with $\xi\in\{S,D\}$ are alpha values. Given the outputs of each branch, we can also render static and dynamic maps separately.
@@ 3.2 Enabling Semantic Awareness
NeRF provides low-level implicit representations of geometry and radiance but lacks a higher-level understanding of the scene. It is shown in [34] that relying only on view-dependent radiance and sensitive mean square error makes the model prone to outliers and over-fit observations. On the other hand, semantic labels can be conceptualized as a view-invariant function [59], mapping world coordinates $\mathbf{x}$ to a distribution over semantic labels $\mathbf{s}(\mathbf{x})\in\mathbb{R}^{L}$, where $L$ is the number of classes. We choose to integrate a semantic head to predict the semantic class for each point, similar to the color head. Subsequently, the same rendering formula is applied for semantics to achieve 2D semantic segmentation maps:
@@
Figure 3: (a) We enforce a “foreground-only mask” to the dynamic semantic head. In the forward pass (black arrows), the mask prevents the dynamic branch from outputting pixels that do not belong to foreground classes. While in the backward pass (orange arrows), it restricts the dynamic field from learning background pixels, which may primarily come from noisy annotations. (b) Predictions generated by 2D segmentation model are noisy and inconsistent between views. (c) Since the dynamic field includes a temporal dimension, it ends up learning these conflicting labels, satisfying the overall loss. (d) Therefore, our proposed mask is used to tackle the problem.
@@
We apply softmax normalization at each query point individually before ray integration, as introduced in [39]. This is done due to the unbounded nature of logits, ensuring that the semantic head can not generate excessively high scores in regions with low density to overcome high-density areas. Remarkably, this not only improves the stability of rendering semantics but also prevents half-precision floating point (float16) overflowing when working with frameworks such as tiny-cuda-nn [24]. By performing inferences in 3D space and rendering to 2D with a foreground masking mechanism, our semantic outputs are multi-view consistent by design.
3.3 Loss Formulation
Robust loss. Our objective is to reduce the sensitivity of the static model on dynamic components through a robust photometric loss. This loss can be formulated as an Iteratively Reweighted Least Squares (IRLS) problem:
where $\epsilon(\mathbf{r})=\|\hat{C}^{S}(\mathbf{r})-C(\mathbf{r})\|_{2}$ is the error residuals and $\omega(\epsilon)=\epsilon^{-1}\cdot\partial\kappa(\epsilon)/\partial\epsilon$ is the weight function of the residuals at the previous iteration. Among the family of kernels [1], we follow Sabour et al. [34] and choose a trimmed estimator [6] which classifies outliers based on a percentile threshold $\mathcal{T}_{\epsilon}$ calculated from the whole batch of rays. This choice ensures outliers are completely removed (i.e. hard weights). Furthermore, as every object has neighbor connections, we apply a spatial filter $\mathcal{F}$ on the residuals map to capture the smoothness, ensuring local support of the kernel:
(a) Ground truth